A jumbo jet taxiing down the runway receives word that it must return to the gate to pick up an important passenger who was late to his connecting flight. The jet is traveling at 45.0 m/s when the pilot receives the message. What is the acceleration of the plane if it takes the pilot 5.00 s to bring the plane to halt?


1
Expert's answer
2021-09-09T18:07:00-0400

Acceleration is defined as the ratio between the change in velocity and the time taken for the velocity to change. Mathematically:

"a=\frac{v-u}{t}"

where

v is the final velocity

u is the initial velocity

t is the time interval

In this problem, we know the following data:

"u=45.0m\/s"  is the initial velocity of the plane

"v=0"  is the final velocity (the plane comes to a stop)

 "t=5.0s" is the time interval

Substituting into the equation, we find

"a=\frac{0-45}{5}" "=-9m\/s" 2

and the negative sign means the acceleration is opposite to the direction of motion of the plane.